OpenStudy (anonymous):

I'm trying to find the field lines, x, of F(x,y,z)=(y-z,-x,x-y). Am I supposed to use the formula:\[\dot{\underline{x}}=\lambda(\underline{x})\underline{F}(\underline{x})\] for some scalarfield lambda? How does it work?

OpenStudy (anonymous):

oh it's F(x,y,z)=(y-z,z-x,x-y)

OpenStudy (anonymous):

\[\dot{\underline{x}}=\frac{d\underline{x}}{dt}\]by the way

OpenStudy (anonymous):

Seems like some kind of method for graphing curl or divergence, something like that but I don't recognize the formula...

OpenStudy (amistre64):

i think it needs more clarification ...

OpenStudy (anonymous):

OK, see if he comes back....

OpenStudy (amistre64):

looks like some sort of differential equation exercise perhaps ...

OpenStudy (anonymous):

Field lines sounds like some sort of contour graph...

OpenStudy (amistre64):

or force field jargon ... the closest thing I can come to online is: http://www.physicsforums.com/showthread.php?t=304499

OpenStudy (anonymous):

Ok, that's a set of calculations, normally interpreting curl,div etc graphically is something of a pain, I would be quite interested if there was some better way of doing that....

OpenStudy (anonymous):

Of course I should have checked Wiki, doesn't explain the formula though... This article is about the modern use of "field lines" as a way to depict electromagnetic and other vector fields. "Modern" , that's why I don't know it, lol.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Fieldline Forgot the link...

OpenStudy (anonymous):

\[\dot{\underline{r}}(t)=(\frac{dx(t)}{dt},\frac{dy(t)}{dt},\frac{dz(t)}{dt})\] So you get a system of linear equations. lambda(x) doesn't have an effect on the direction so you can choose anything you want for that I believe. The answer to this question is supposedly \[x^{2}+y^{2}+z^{2}=c_{1} \]\[x+y+z=c_{2}\] so you get circles.

OpenStudy (anonymous):

The line underneath letters means that they're vectors.

OpenStudy (anonymous):

Yes, I got that, I will have another look at this a bit later on...

OpenStudy (anonymous):

Did I say linear equations, I meant differential equations.